OpenStudy (anonymous):

among all the pairs of numbers whose difference is 16 find a pair whose product is as small as possible. what is the minimum product?

OpenStudy (anonymous):

maybe -8 and 8 whose product is -64 just a guess though

OpenStudy (anonymous):

x-y=16 x=16+y xy=y(16+y)=16y+y^2 derive: 2y+16=0 y=-8 x=8

OpenStudy (amistre64):

|small| or -small?

OpenStudy (anonymous):

\[x-y=16\] \[x=16+y\] \[f(x)=xy=x(16+x)\] now derivative and etc

OpenStudy (anonymous):

smallest product -64

OpenStudy (anonymous):

y - x = 16 y = x + 16 x(x+16) = x^2 + 16x vertex is at 8 x = -8 y = 8 minimum product is -64

OpenStudy (anonymous):

it just says small

OpenStudy (anonymous):

minimum means smallest

OpenStudy (anonymous):

i would argue from symmetry, not from calculus. x - y = 16 or y - x = 16 totally symmetric